OpenStudy (ttp):

The length of the hypotenuse of a right triangle is eighteen centimeters and the length of one side is twelve centimeters. What is the exact length of the third side? square root of six centimeters square root of one hundred eighty centimeters square root of three hundred twelve centimeters square root of four hundred sixty-eight centimeters

OpenStudy (calculusxy):

Use the Pyhtagorean Theorem: a^2+b^2=c^2 12+b^2=18 Solve for b^2.

OpenStudy (anonymous):

"square root of one hundred eighty centimeters"\[\sqrt{18^2-12^2}=\sqrt{180} \]

OpenStudy (the_fizicx99):

o.0 That would be 12^2 + b^2 = 18^2 144 + b^2 = 324 -144 -144 b^2 = 180 Square root

OpenStudy (anonymous):

Thank you for the medal TTP

OpenStudy (ttp):

Thanks! @robtobey @tHe_FiZiCx99 @calculusxy

OpenStudy (ttp):

Welcome! @robtobey

OpenStudy (ttp):

I would give @tHe_FiZiCx99 but i can't can somebody? :)

OpenStudy (the_fizicx99):

Wot?

OpenStudy (ttp):

i wanna give y'all boh a metal haa